# Collaborating with AI Agents: Field Experiments on Teamwork, Productivity, and Performance

### v1

@@ 2.4 Participants
We pre-registered a sample of 2,500 participants from Prolific (www.prolific.com) based in the US, with representative stratification across gender and ethnicity. In total, 2310 participants completed the task. Of the 2500 participants who entered, 23 did not enter the matching queue (see Section 2.1 for details) and were removed from the study. A further 167 participants quit the study or were timed out before they were matched to a partner. The overall attrition rate was 7.6%. A balance check found that our sample was balanced on all measured covariates (Table 1). This study was run from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid 9 US dollars, and two participants were each awarded a bonus payment of \$100 who won a lottery based on productivity and performance.
@@ 2.5 Summary statistics and randomization checks
In total, the dataset includes 2,310 participants, 1,834 teams, 11,138 display ad submissions, 183,691 messages, 63,656 edits on images, 1,960,095 edits on ad copy, and 10,375 AI-generated images. To ensure that the randomization procedure successfully balanced covariates across experimental conditions, Table 1 compares key participant characteristics between the Human-Human and Human-AI conditions. These include demographic variables (e.g., gender, age) and psychological traits based on the Big Five personality dimensions. As shown, no significant differences were detected, indicating successful randomization.
Covariate All Human-AI Human-Human
Individuals 2310 1258 1052
Teams 1834 1258 576
Gender (% Male) 50.8% 50.4% 51.3%
Age 42 $\pm$ 14 43 $\pm$ 15 42 $\pm$ 14
Openness 0.71 $\pm$ 0.20 0.71 $\pm$ 0.21 0.72 $\pm$ 0.19
Conscientiousness 0.79 $\pm$ 0.18 0.78 $\pm$ 0.18 0.79 $\pm$ 0.17
Extraversion 0.56 $\pm$ 0.21 0.55 $\pm$ 0.21 0.56 $\pm$ 0.21
Agreeableness 0.70 $\pm$ 0.22 0.70 $\pm$ 0.21 0.69 $\pm$ 0.24
Neuroticism 0.48 $\pm$ 0.23 0.49 $\pm$ 0.24 0.48 $\pm$ 0.22
Table 1: Randomization check. Personality traits are normalized from a 7-point Likert scale.
@@ 2.6 Incentives
To incentivize participants to create high-quality ads, we informed them that eligibility for an additional \$100 prizes would be based on both the quality and quantity of their ads, as well as their performance on social media platforms. Participants were explicitly instructed that "the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each for producing the best-performing ads.
@@ 2.7 Message labeling
To label messages into categories and intent, we prompt gpt-4o-mini-2024-07-18 for each message independently and ask for a label for a category and an intent of the message. We enforce that the labels are from the set of pre-determined labels using OpenAI’s Structured Outputs API. See Appendix A for the prompts used in this analysis.
@@ 2.8 AI evaluation of ad quality
To obtain ratings of display ads that are as close as possible to display ads one would see on digital ad publishers, we first created mockups of each display ad. We built a web application that populates the image, ad copy, shortened link, call-to-action, and other user interface items—including the Like, Comment, Share, and Close buttons, the profile picture, and the Sponsored tag. Screenshots were then programmatically taken of each mockup. See Figure 3 for examples.
@@
If AI ratings are shown to predict human evaluations of ad quality and field evaluations of ad performance, they can potentially provide cost-effective alternatives to ad testing. To obtain AI ratings of ad quality, we prompt OpenAI’s multimodal gpt-4o-mini-2024-07-18—which supports image input and structured outputs. [^3] See OpenAI’s documentation on vision and structured outputs. To make AI evaluations comparable to human evaluations, we ask the same questions as those given to human evaluations (see Section 2.9). Each item was on a 7-point Likert scale. The first question was "The text is present, clear, relevant, and engaging"; the second "The image is visually appealing"; and the third "I am likely to click on this ad." The exact prompts are shown in Appendix C.
@@ 2.9 Human evaluation of ad quality
To obtain human ratings of ad quality, we recruited a separate sample of 1,995 participants from Prolific based in the US, with representative stratification across gender, age, and ethnicity. A total of 1,200 individuals entered the survey. Of these participants, 5 dropped out before submitting their surveys, resulting in a dropout rate of 0.42%. We used our custom platform for this survey, run on Google Cloud Platform’s App Engine. The code is available on GitHub. This study was not a part of our pre-registration. This study was run from November 7 to 9, 2024, and took a median of 16.47 minutes to complete.
@@
We used the same mockups of display ads as used for AI evaluation of ad quality, as explained in Section 2.8. For each display ad, we asked the participants three questions regarding the quality of the text, the image, and the estimated clickthrough rate. Each item was on a 7-point Likert scale. The first question was "The text is present, clear, relevant, and engaging"; the second "The image is visually appealing"; and the third "I am likely to click on this ad." See Figure 4 for an example of an item on the survey.
@@ 2.10 Field evaluation of ad performance
To assess real-world advertising outcomes, we ran ad campaigns on X (formerly Twitter), pre-registered on osf.io/95dhu. We use the following as outcome variables: click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; as a fraction of document viewed), and view-through duration (VTD; in seconds). We tracked VTR and VTD by using a unique link on DocSend for each ad.
Due to the effects of divergent targeting in A-B testing (Braun and Schwartz, 2024), we follow recommendation 5 from Braun et al. (2024) and run a multi-ad study to test the causal evidence of the impact of ads conditional on online algorithms. We do not use holdout in this study; each ad has a unique DocSend link so we expect the outcome to be zero in a holdout group and will attribute all document visits to the ads.
To run as as many ads as possible while sampling evenly from human-human and human-AI teams and from the human predictions of ad quality, we implemented a stratified sampling approach. We obtained a sample of 2,000 ads from a total of 11,138 ads, sampling between one and two ads from each of the 1,834 teams. We first divided the ads into those created by human-human teams and human-AI teams and then further divided the ads into 10 strata according to the click scores, ranging from low to high predicted click probability. The click scores were the human predictions of the likelihood of user clicks (see Section 2.9). We removed 8 ads that potentially violated moderation policies (e.g., violence, sexual, drug) from the total set. We controlled for spend in our analyses to account for auto-bidding, which we had left on for our campaigns.
To prevent overlapping audience targeting, we assigned five unique ads to each of 400 campaigns, structuring them as 5-ad split tests within individual campaigns (noting the platform’s limit of five split tests per campaign). [^4] The platform limit for split tests within each campaign is 5. To further ensure no overlap across campaigns, we allocated a random set of 133 ZIP codes to each campaign, selecting ZIP codes with populations between 10,000 and 100,000.555The data is from https://data.census.gov/table/ACSDT5Y2020.B01003. We tested the robustness of this allocation across ZIP codes by conducting one-way ANOVA on population and income, yielding non-significant results: $F(399,53199)=0.954$, $p=0.734$ for population, and $F(399,53199)=0.973$, $p=0.636$ for income.
@@ 2.11 Model specifications
We measure the effects of human-AI collaboration using a standard regression model. For individual $i$, we use the following model:
$Y_{i}=\delta\text{H-AI}_{i}+\sum_{p\in\text{Big Five}}\beta_{p}p_{t}+\beta_{1}\text{Age}_{i}+\beta_{2}\text{Gender}_{i}$ (1)
where $Y_{t}$ represents the outcome for team $t$, H-AIi is 1 if individual $i$ is working with an AI and 0 if individual $i$ is working with another participant, $p$ is a normalized score for each Big Five trait, and Genderi is 1 if the participant is male and 0 if female. The outcomes of interest $Y_{t}$ include the number of interactions on the platform (i.e., messages, copy edits, image selects, AI image generations, submissions), fraction of messages by category, quality scores, and fraction of copy completed. For the analysis of ad quality, we perform analysis on the rating level where $i$ is an ad rating.
For team-level analysis for team $t$, we use the following model:
$Y_{t}=\delta_{t}\text{H-AI}_{t}+\sum_{p\in\text{Big Five}}\beta_{p}\bar{p}_{t}+\sum_{i\in\text{Team}_{t}}\beta_{1i}\text{Age}_{ti}++\sum_{g\in\{M,F\}}\text{Gender}_{g}\text{,}$ (2)
where the outcome $Y_{t}$ is the same as above, H-AIt (Human-AI) is 1 if the team was Human-AI and 0 if the team was Human-Human, $\bar{p}_{t}$ is the average Big Five trait for both participants, and Genderg is 1 if both are $g\in\{M,F\}$ and 0 if one participant is male and the other is female.
Prompt effects.
[…]
$Y_{i}$ $=\sum_{P_{AI}\in\text{Big Five}}P_{AI}+\sum_{P_{H}\in\text{Big Five}}\beta_{P_{H}}P_{H,i}+\beta_{1}\text{Age}_{i}+\beta_{2}\text{Gender}_{i}$
$+\sum_{P_{AI}\in\text{Big Five}}P_{AI}\times\Bigg(\sum_{P_{H}\in\text{Big Five}}\beta_{P_{AI},P_{H}}P_{H,i}+\beta_{P_{AI},1}\text{Age}_{i}+\beta_{P_{AI},2}\text{Gender}_{i}\Bigg)$
where the outcome $Y_{t}$ is the same as above, $P_{AI}$ is the orthogonally randomized, prompted high (i.e., 1) or low (i.e., 0) AI personality (Jiang et al., 2023).
@@
$Y_{tc}=\delta_{t}\text{H-AI}_{tc}+\text{Image}_{tc}+\text{Text}_{tc}+\text{Click}_{tc}+\text{Spend}_{tc}+u_{c}\text{,}$ (3)
@@ 3 Results
3.1 Teamwork
Collaborating with AI increases communication but decreases direct copy edits.
To test how human-AI collaboration affects communication, we measured the number of chat messages sent by participants in both Human-Human and Human-AI groups. In the study, participants worked collaboratively with and sent messages to their partners in real-time, for both Human-Human and Human-AI groups. In the Human-Human group, participants messaged their human partners; in the Human-AI group, participants messaged the AI as one does with chatbots. The MindMeld platform logs a timestamped record of each message sent by a participant or AI.
In our study, participants communicated significantly more in Human-AI teams than in Human-Human teams. Participants in Human-AI teams sent 45% more messages than those in Human-Human teams (Table 2). Interestingly, personality covariates were significantly associated with the number of messages sent in Human-AI teams: individuals scoring higher on openness and extraversion sent more messages, while those scoring higher on agreeableness sent fewer messages. These results indicate that collaborating with an AI partner encourages more frequent communication and that individual differences in personality traits, such as openness, extraversion, and agreeableness, further influence communication patterns in these interactions.
@@ 3.1 Teamwork
On the other hand, participants in Human-AI teams made significantly fewer direct copy edits compared to Human-Human teams, with a 60% decrease in edits (Table 3. The results on copy edits and messages suggest that collaboration with AI shifted participants’ focus away from iterative textual refinements to edits through instructions and suggestions to the AI, as we will see further in the next analyses. A post hoc analysis of the correlation between user actions found that the number of messages sent and copy edits were negatively correlated ($R^{2}=-0.17$, $p=3.5^{-17}$), further supporting this shift. However, participants in the Human-AI condition still engaged in direct editing, indicating that while AI collaboration reduced the frequency of manual edits, it did not eliminate them entirely. Unlike previous studies, such as Chen and Chan (2024), that artificially constrained participants to specific modalities when interacting with AI, our results show that in more realistic collaborative settings with AI agents, participants naturally utilized both direct editing and interaction through messaging.
@@
In addition to the number of messages sent across conditions, we investigate whether the types of messages sent vary across conditions. We used gpt-4o-mini to label each message independently with one of five message categories: Process, Content, Social, Emotional, Feedback, or Other (see Section 2.7 for more details). In our analysis, we found that both the AI and the individuals in Human-AI teams sent more content- and process-oriented messages, while Human-Human teams sent more social and emotional messages (Figure 5). This shift indicates that collaboration with AI emphasizes task-related communication over social interaction, possibly because participants can focus more on the task without needing to navigate the social or emotional aspects of collaboration.
@@
To further examine how communication differs across conditions, we labeled the messages with one of 36 message intents (see Appendix A for all labels). These intents capture a range of functions, such as suggestions, instructions, and prioritization, as well as rapport building and self-assessment. We found notable differences in the distribution of message intents between the Human-Human and Human-AI groups in each message category (Table 4). See our model-free evidence in Figure 6. For example, in the Content category, Human-Human teams send more messages that show confusion and clarification, and Human-AI teams send more messages about brainstorming, confirmation, acknowledgment, suggestion, agreement, instruction, and judgment. The results are similar for Process-related messages. For messages in the Social and Emotional categories, Human-Human teams send more messages about humor, concern, and apologies, while Human-AI teams send more messages about suggestions, appreciation, motivation, confirmation, and satisfaction. These findings further support the idea that Human-AI collaboration prioritizes task-oriented communication, while Human-Human teams engage more in social and emotional exchanges.
@@
Survey reports showed no differences in teamwork.

### v2

@@ 2.4 Participants
We pre-registered a sample of 2,500 participants from Prolific (www.prolific.com) based in the US, with representative stratification across gender and ethnicity. In total, 2310 participants completed the task. Of the 2500 participants who entered, 23 did not enter the matching queue (see Section 2.3 for details) and were removed from the study. A further 167 participants quit the study or were timed out before they were matched to a partner. The overall attrition rate was 7.6%. The study was conducted from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid \$9.
@@ 2.5 Summary statistics and randomization checks
In total, the dataset includes 2,310 participants, 1,834 teams, 11,138 display ad submissions, 183,691 messages, 63,656 edits on images, 1,960,095 edits on ad copy, and 10,375 AI-generated images. To ensure that the randomization procedure successfully balanced covariates across experimental conditions, Table 1 compares key participant characteristics between the human-human and human-AI conditions. These include demographic variables (e.g. age, gender), participants’ work status, and the Big Five personality traits. No significant differences were detected apart from a very small difference in part-time employment status across treatment conditions, indicating broad-based, balanced randomization across participants.
Variables All Human-AI Human-Human t-statistic (SE) p-value
Individuals 2310 1258 1052 - -
Teams 1834 1258 576 - -
Gender (% Male) 50.8% 50.4% 51.3% -0.401 (0.021) 0.688
Age 42 $\pm$ 14 43 $\pm$ 15 42 $\pm$ 14 0.987 (0.599) 0.324
Full-Time 47.8% 48.5% 47.1% 0.688 (0.021) 0.491
Part-Time 14.7% 13.4% 16.3% -2.009 (0.015) 0.045
Data expired 13.4% 13.2% 13.6% -0.279 (0.014) 0.780
Not in paid work 10.3% 10.9% 9.7% 0.943 (0.013) 0.346
Unemployed (and job seeking) 8.8% 9.1% 8.6% 0.428 (0.012) 0.668
Other 4.2% 4.2% 4.3% -0.077 (0.008) 0.939
Start new job in a month 0.6% 0.8% 0.5% 0.974 (0.003) 0.330
Openness 0.71 $\pm$ 0.20 0.71 $\pm$ 0.21 0.72 $\pm$ 0.19 -0.854 (0.008) 0.393
Conscientiousness 0.79 $\pm$ 0.18 0.78 $\pm$ 0.18 0.79 $\pm$ 0.17 -1.166 (0.007) 0.244
Extraversion 0.56 $\pm$ 0.21 0.55 $\pm$ 0.21 0.56 $\pm$ 0.21 -0.962 (0.009) 0.336
Agreeableness 0.70 $\pm$ 0.22 0.70 $\pm$ 0.21 0.69 $\pm$ 0.24 0.534 (0.009) 0.594
Neuroticism 0.48 $\pm$ 0.23 0.49 $\pm$ 0.24 0.48 $\pm$ 0.22 1.538 (0.010) 0.124
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Personality traits are normalized from a 7-point Likert scale. For t-statistics, gender is coded as 1 for male and 0 for female, and employment status is 1 for the listed category (e.g., full-time) and 0 otherwise.
Table 1: Randomization checks
@@ 2.6 Incentives
To incentivize participants to create high-quality ads, we informed them that they were eligible for additional \$100 prizes based on the quantity and quality of the ads they submitted, as well as the performance of the ads in the marketing field experiment. Participants were explicitly instructed that “the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each for producing the greatest number of high-performing ads.
@@ 2.7 Message labeling
To label messages into categories and intent, we prompt gpt-4o-mini-2024-07-18 for each message independently and ask for a label for a category and an intent of the message. We enforce that the labels are from the set of pre-determined labels using OpenAI’s Structured Outputs API.
@@ 2.8 AI evaluation of ad quality
To obtain ratings of display ads that are as close as possible to display ads one would see on digital ad publishers, we first created mockups of each display ad. An ad mockup is a simulated representation of a display ad designed to closely resemble the appearance and functionality of an actual ad as it would appear on digital ad publishers’ platforms. We built a web application that populates the image, ad copy, shortened link, call-to-action, and other user interface items—including the Like, Comment, Share, and Close buttons, the profile picture, and the Sponsored tag. Screenshots were then programmatically taken of each mockup. See Figure 3 for examples.
@@
If AI ratings are shown to predict human evaluations of ad quality and field evaluations of ad performance, they can potentially provide cost-effective alternatives to ad testing. To obtain AI ratings of ad quality, we prompt OpenAI’s multimodal gpt-4o-mini-2024-07-18, which supports image input and structured outputs. [^3] See OpenAI’s documentation on vision and structured outputs. To make AI evaluations comparable to human evaluations, we asked the same questions of the AI as those given to human evaluators (see Section 2.9). Each item was evaluated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad."
@@ 2.9 Human evaluation of ad quality
To obtain human ratings of ad quality, we recruited a separate sample of 1,995 participants from Prolific based in the US, with representative stratification across gender, age, and ethnicity. A total of 1,200 individuals entered the survey. Of these participants, 5 dropped out before submitting their surveys, resulting in a dropout rate of 0.42%. We built a custom platform for this survey, run on Google Cloud Platform’s App Engine. [^4] The code is available on GitHub. This evaluation was conducted from November 7 to 9, 2024, and took a median of 16.47 minutes to complete.
@@
We used the same mockups of display ads as used for AI evaluation of ad quality, as explained in Section 2.8. For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." See Figure 4 for an example of an item on the survey.
@@ 2.10 Field evaluation of ad performance
To assess real-world advertising outcomes, we ran ad campaigns on the social media platform X, pre-registered on osf.io/95dhu. We use the following as outcome variables: click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; as a fraction of document viewed), and view-through duration (VTD; in seconds). We tracked VTR and VTD by using a unique link on DocSend for each ad.
Due to the effects of divergent targeting in A-B testing (Braun and Schwartz, 2024), we follow recommendation 5 from Braun et al. (2024) and run a multi-ad study to test the causal effects of the ads on performance, conditional on online algorithms. We do not use a holdout in this study. Each ad has a unique DocSend link, so we expect the outcome in a hypothetical holdout group to be zero and attribute all document visits to the ads (Braun et al., 2024).
To run as many ads as possible while sampling evenly from human-human and human-AI teams and from the human predictions of ad quality, we implemented a stratified sampling approach. We obtained a sample of 2,000 ads from a total of 11,138 ads, sampling between one and two ads from each of the 1,834 teams. We first divided the ads into those created by human-human teams and human-AI teams and then further divided the ads into 10 strata according to the human predictions of the likelihood of user clicks, from low to high predicted click probability (see Section 2.9). We removed 8 ads that potentially violated moderation policies (due to the inclusion of e.g. violence, sexual, or drug content). We controlled for advertising spend in our analyses to account for auto-bidding.
To prevent overlapping audience targeting, we assigned five unique ads to each of 400 campaigns, structuring them as 5-ad split tests within individual campaigns (noting the platform’s limit of five split tests per campaign). To further ensure no overlap across campaigns, we allocated a random set of 133 ZIP codes to each campaign, selecting ZIP codes with populations between 10,000 and 100,000.555The data is from https://data.census.gov/table/ACSDT5Y2020.B01003. We tested the robustness of this allocation across ZIP codes by conducting one-way ANOVA on population and income, yielding non-significant results: $F(399,53199)=0.954$, $p=0.734$ for population, and $F(399,53199)=0.973$, $p=0.636$ for income.
@@ 2.11 Model specifications
Due to the experimental nature of our design, we measure the effects of human-AI collaboration using a standard regression model:
$Y_{i}=\delta\text{HAI}_{i}+\beta X_{i}+\epsilon_{i}\text{,}$ (1)
where $Y_{t}$ represents the outcome for team $t$, HAIi is 1 if individual $i$ is working with an AI and 0 if individual $i$ is working with another participant, $X_{i}$ includes demographic controls including age, gender, and Big Five personality traits, and $\epsilon_{i}$ is the error term. The outcomes of interest $Y_{t}$ include the number of interactions on the platform (i.e., messages, copy edits, image selects, AI image generations, submissions), fraction of messages by category, quality scores, and fraction of copy completed. For the analysis of ad quality, we perform analysis on the rating level where $i$ is an ad rating. For team-level analysis, we use the following model:
$Y_{t}=\delta_{t}\text{HAI}_{t}+\beta X_{t}+\epsilon_{t}\text{,}$ (2)
where the outcome $Y_{t}$ is the same as above, HAIt (Human-AI) is 1 if the team was Human-AI and 0 if the team was Human-Human, $X_{t}$ includes team-level demographic controls including average age, gender composition, and average Big Five personality traits, and $\epsilon_{t}$ is the error term.
@@
$Y_{tc}=$ $\ \delta_{t}\text{HAI}_{tc}+\text{Text}_{tc}+\text{Image}_{tc}+\text{Click}_{tc}$
$+\text{Spend}_{tc}+u_{c}+\epsilon_{tc}\text{,}$ (3)
@@ 3 Results
3.1 Work Processes
Collaborating with AI increases communication and decreases direct engagement with work product.
To test how human-AI collaboration affects communication, we measured the number of messages sent by participants in both human-human and human-AI groups. Participants worked collaboratively with and sent messages to their partners in real-time, for both human-human and human-AI groups. In the human-human group, participants messaged their human partners; in the human-AI group, participants messaged the AI. The Pairit platform logs a timestamped record of each message sent by a participant or AI.
Participants communicated significantly more in human-AI teams than in human-human teams. Participants in human-AI teams sent 63% more messages than those in human-human teams, with consistent results when demographic covariates are included (Table 2). These results indicate that collaborating with an AI partner encourages more frequent communication.
@@ 3.1 Work Processes
On the other hand, participants in Human-AI teams made 71% fewer direct copy edits compared to Human-Human teams (Table 3). The results on copy edits and messages suggest that collaboration with AI shifted participants’ focus from generating text to instructing and suggesting the AI to generate text, as we will see in greater detail in the next set of analyses. A post hoc analysis of the correlation between user actions found that the number of messages sent and copy edits were negatively correlated ($R^{2}=-0.17$, $p=3.5^{-17}$), further supporting this interpretation. However, participants in the Human-AI condition still engaged in direct editing, indicating that while AI collaboration reduced the frequency of manual edits, it did not eliminate them entirely. Unlike previous studies, such as Chen and Chan (2024), which artificially constrained participants to specific modalities when interacting with AI, our results show that in more realistic collaborative settings, where humans and agents can freely allocate effort to different tasks, participants utilized both direct engagement with work product (ad editing) and interaction through messaging.
@@
In addition to the number of messages sent across conditions, we investigated whether the content of messages varied across conditions. We used gpt-4o-mini to label each message independently with one of five message categories: Process, Content, Social, Emotional, Feedback, and Other (see Section 2.7 for more details). In our analysis, we found that both the AI and the humans in human-AI teams sent 18% (i.e., (0.4060.345)/0.345=0.177(0.406-0.345)/0.345=0.177) more content- and process-oriented messages, while human-human teams sent 29% (i.e., (0.5790.449)/0.449=0.29(0.579-0.449)/0.449=0.29) more social and emotional messages (Figure 5, Table 4). This shift indicates that collaboration with AI emphasizes task-related communication over social interaction, possibly because participants can focus more on the task without needing to navigate the social or emotional aspects of collaboration.
@@
To further examine how communication differs across conditions, we labeled the messages with one of 36 message intentions. These intentions capture a range of functions, such as suggestions, instructions, and prioritization, as well as rapport building and self-assessment. We found notable differences in the distribution of message intentions between the human-human and human-AI groups in each message category (see Table 4 and Figure 6). For example, in the Content category, human-human teams send more messages that exhibit confusion and clarification, and human-AI teams send more messages about brainstorming, confirmation, acknowledgment, suggestion, agreement, instruction, and judgment. The results are similar for process-related messages. For messages in the social and emotional categories, human-human teams send more apologies and messages about humor and concern, while human-AI teams send more suggestions and messages about appreciation, motivation, confirmation, and satisfaction. These findings further support the idea that human-AI collaboration prioritizes task-oriented communication, while human-human teams engage more in social and emotional exchanges.
@@
Survey reports showed no differences in perceptions of teamwork quality.